OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) - 4 I'm not looking for someone to tell me the answer, so please do not post it. I'm looking for someone to help me figure out how to answer this question :/

OpenStudy (anonymous):

alright, distribut the 3 to the x and the 1; then distribute the -2 to the x and the -1; that would give you 3x+3 = -2x +2 - 4, add 2x to both sides, subtract 3 from both sides, nd that would give you 5x=5-4; so that would give you 5x=1, divide 5 from both sides and that gives you the answer

OpenStudy (anonymous):

So the answer would be 1?

OpenStudy (anonymous):

@savannaxx_

OpenStudy (anonymous):

no the answer would be 1/5